OpenStudy (anonymous):

3x^2y^2 - 75 can someone help me they said to also use F.O.I.L

OpenStudy (anonymous):

i don't know who the "they" is, but they are dead wrong first off, there is no such thing as foil secondly your job it to factor, not multiply out

OpenStudy (anonymous):

start with \[3(x^2y^2-25)\] then factor the \(x^2y^2-25\) part as the difference of two squares \[a^2-b^2=(a+b)(a-b)\] with \(a=xy,b=5\)

OpenStudy (anonymous):

they told me i could use foil but plzzz can you help me :(

OpenStudy (anonymous):

they are full of crap

OpenStudy (anonymous):

lol ikw right

OpenStudy (anonymous):

did you understand what i wrote above?

OpenStudy (anonymous):

kinda im sorry im just trying to finish this thing they keep pressuring me about it and i dont understand it sorry

OpenStudy (anonymous):

start with \[3x^2y^2-75\] both terms have a common factor of \(3\) so "factor it out" and get \[3(x^2y^2-25)\]

OpenStudy (anonymous):

then factor the \((x^2y^2-25)\) piece as the difference of two squares \((a+b)(a-b)\) with \(a=xy, b=5\)

OpenStudy (anonymous):

this will give you \[3(xy+5)(xy-5)\]

OpenStudy (anonymous):

is there really a "they"?

OpenStudy (anonymous):

yeah my online teachers

OpenStudy (anonymous):

if your on line teacher said "foil" they should be banned from teaching

OpenStudy (anonymous):

can you help me with more plzzz :/ ill give u medals i promise

OpenStudy (anonymous):

and this is not multiplying, it is factoring

OpenStudy (anonymous):

lol right i dont thing they know wat they doing

OpenStudy (anonymous):

i will be happy to help as for the medals, i have plenty

OpenStudy (anonymous):

i know they don't

OpenStudy (anonymous):

lol okay but i would love to make u have more its not a bother to me

OpenStudy (anonymous):

go ahead and ask

OpenStudy (anonymous):

ok 2x^3 + 54

OpenStudy (anonymous):

factor? first factor out the 2 and get \[2(x^3+27)\] then you have the "sum of two cubes" which factors as \[a^3+b^3=(a-b)(a^2-ab+b^2)\] in your case \(a=x,b=3\) giving you \[2(x+3)(x^2-3x+9)\]

OpenStudy (anonymous):

sorry middle line is wrong, it is \(a^3+b^3=(a+b)(a^2-ab+b^2)\) answer is correct though

OpenStudy (anonymous):

oh okay lol but i have more if you dont mind :/ im sorry

OpenStudy (anonymous):

you have me for ten more minutes go ahead and ask

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

3. 2x^2 - 11x - 40

OpenStudy (anonymous):

ok i hate factoring, but we can do it

OpenStudy (anonymous):

okay:)

OpenStudy (anonymous):

i hope you don't have to explain how to do it, do you? answer is \[(x-8)(2x+5)\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok then let me show you how to cheat http://www.wolframalpha.com/input/?i=2x^2+-+11x+-+40+

OpenStudy (anonymous):

next...

OpenStudy (anonymous):

5. x^2 – x – 42

OpenStudy (anonymous):

omg they give u everything thankkkk you :O

OpenStudy (anonymous):

this one is easier since \(-7\times 6=-42\) and \(-7+6=-1\) your answer is \[(x-7)(x+6)\]

OpenStudy (anonymous):

lol easy right?

OpenStudy (anonymous):

lol yeah

OpenStudy (anonymous):

of course it doesn't explain much any more?

OpenStudy (anonymous):

really

OpenStudy (anonymous):

2x^2 – 20x + 40

OpenStudy (anonymous):

factor out the 2 and get \[2(x^2-10x+20)\]

OpenStudy (anonymous):

that is the best you can do, because \(x^2-10x+20\) does not factor using integers

OpenStudy (anonymous):

yeah i kinda saw that 2 lol

OpenStudy (anonymous):

that is really all you can do for this one

OpenStudy (anonymous):

any more?

OpenStudy (anonymous):

6. 21x^2 + 31x + 4

OpenStudy (anonymous):

7. 3x^3 + 27

OpenStudy (anonymous):

\[(3 x+4) (7 x+1)\]for the first one

OpenStudy (anonymous):

8. x^2 + 9x + 20

OpenStudy (anonymous):

second one it is really \(3x^3+27\)? only can factor out a 3 and get \(3(x^3+9)\)

OpenStudy (anonymous):

as for number 8, see if you can do it yourself if you cannot, paste it in to wolfram and see what you get

OpenStudy (anonymous):

0h okay np

OpenStudy (anonymous):

for \(x^2 + 9x + 20\) see if you can think up two numbers that if you multiply them you get 20, and if you add them you get 9 it should be pretty easy here

OpenStudy (anonymous):

with a tiny bit of practice you will come up with \(4\) and \(5\)right away, since \(4\times 5=20\) adn \(4+5=9\)

OpenStudy (anonymous):

this means \[x^2+9x+20=(x+4)(x+5)\]

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

any more?

OpenStudy (anonymous):

yeah i have one last one but it looks diffrent

OpenStudy (anonymous):

okay lets take a look?

OpenStudy (anonymous):

3rs - 2r - 3s + 2

OpenStudy (anonymous):

summer on line class?

OpenStudy (anonymous):

yeah its killing me

OpenStudy (anonymous):

must be punishment for failing in the spring huh?

OpenStudy (anonymous):

yeah i had a really crappy teacher and i didnt learn much

OpenStudy (anonymous):

\[3rs - 2r - 3s + 2\]\[=r(3x-1)-1(3s-2)=(r-1)(3s-2)\] this is called "factor by grouping" mostly i don't believe it when people say they failed because of a crappy teacher, but if your teacher say "foil" i believe you good luck! done?

OpenStudy (anonymous):

awww thanks :) and yes i am thank you sooo much ill message u wen i need anything else :)

OpenStudy (anonymous):

you can always come back and post when you need help

OpenStudy (anonymous):

ok :)